# A Fast, Analytic Empirical Model of the Gaia Data Release 3 Astrometric Orbit Catalog Selection Function

### v1

@@ 5.2 Probability of finding Gaia black holes
Note. — Properties of Gaia BH1, BH2, and BH3. * subscript denotes properties of the stellar companion, $\bullet$ subscript denotes properties of the black hole. All values except $M_{G,*}$ and $A_{G}$ are taken from El-Badry et al. (2023a), El-Badry et al. (2023b), and Gaia Collaboration et al. (2024) for Gaia BH1, BH2, and BH3, respectively. We note for Gaia BH2 that our extinction map suggests much less extinction ($\approx 0.4$ mag) than the map used by El-Badry et al. (2023b) and so our inferred extinction-corrected absolute magnitude of the star $M_{G,0*}$ in Gaia BH2 is fainter. Similarly for Gaia BH3 our extinction map suggests $\approx 0.3$ mag less extinction than Gaia Collaboration et al. (2024).
@@
We apply our forward modeling framework to determine the probability of inclusion of Gaia BH1 (El-Badry et al., 2023a; Chakrabarti et al., 2023), Gaia BH2 (El-Badry et al., 2023b), and Gaia BH3 (Gaia Collaboration et al., 2024) in the Gaia DR3 NSS astrometric orbits catalog. We calculate the probability as a function of sky position, orbital period, eccentricity, parallax, and the component magnitudes and masses. We integrate over random inclinations in $|\cos i|$.
@@
We find the probability of Gaia BH1 ((RA, Dec) = (262.17, –0.58) deg, $P_{orb}=186$ day, $e=0.45$, $\varpi=2.09$ mas, $M_{*}=0.93M_{\odot}$, $M_{\bullet}=9.62M_{\odot}$, $G_{*}=13.77$ mag) being included in the Gaia DR3 NSS astrometric orbit catalog to be 33%. We also vary the location of BH1 on the sky and its distance from the Sun (Figure 9) and consider how its probability of inclusion in the Gaia DR3 NSS astrometric orbit catalog changes. Since Gaia BH1 was located near the ecliptic plane, it had a relatively low probability of inclusion; toward the ecliptic poles, the inclusion probability would have been closer to 70–90%. Gaia BH1 was also relatively close by, at a distance of about 500 pc. Had it been at the same (RA, Dec) but slightly farther away, at 600 pc or more, it would not have made it into the NSS catalog. Conversely, had it been slightly closer, 400 pc or less, it would have had $>80\%$ probability of making it into the catalog. This suggests that if there were Gaia BH1-like systems within a few hundred pc of the Sun, they should have been included in the Gaia DR3 NSS astrometric orbits catalog; the lack of such detections suggests that there are no or few Gaia BH1-like sources within 300–400 pc of the Sun.
@@
We find the probability of Gaia BH2 ((RA, Dec) = (207.57, –59.24), $P_{orb}=1277$ day, $e=0.52$, $\varpi=0.86$ mas, $M_{*}=1.07M_{\odot}$, $M_{\bullet}=8.94M_{\odot}$, $G_{*}=12.28$ mag) being included in the Gaia DR3 NSS astrometric orbit catalog to be 4%. We also vary the location of BH2 on the sky and its distance from the Sun (Figure 9) and consider how its probability of inclusion in the Gaia DR3 NSS astrometric orbit catalog varies. Since Gaia BH2 has a long orbital period, the main factor of whether it made it into the NSS catalog is whether or not it made it through the astrometric cascade or not. The inclusion probability of Gaia BH2 does not change significantly with distance until beyond $\gtrsim 2.5$ kpc, at which point the inclusion probability substantially drops. The inclusion probability also does not vary significantly across the sky. This suggests that the main factor of whether Gaia BH2 made it through the cascade or not is the long period filtering. The probability of an orbital solution being kept at $P_{orb}=1277$ days is 4% (Figure 5). This suggests that Gaia BH2 was a relatively lucky detection and that there are many more such systems that will be detectable in Gaia DR4, either in an updated NSS catalog, or in the epoch astrometry.
We find the probability of Gaia BH3 to be included in the Gaia DR3 NSS astrometric catalog to be 0%. This is because of the orbital period being $>4000$ days; none of these systems pass the long period filtering. This is not entirely correct, as $<0.5\%$ of solutions in the NSS astrometric orbits catalog have $P_{orb}>2000$ days, and $<0.05\%$ of solutions have $P_{orb}>4000$ days, so it is possible for such long solutions to be in the catalog. However, within a margin of a few percent, this is probably accurate. In addition, Gaia BH3 was not included in any of the NSS catalogs, orbital (nss_two_body_orbit), nor acceleration (nss_acceleration_astro, nss_non_linear_spectro).
@@
Figure 9: Top panels: Probability of inclusion of Gaia BH1 (left) and Gaia BH2 (right) in Gaia DR3 NSS astrometric orbits catalog as a function of sky position (192 equally spaced points on sky); the color of the circle denotes the probability of inclusion. Note that the color bar scales in the two panels are different. We fix all parameters as listed in Table 2, and only vary the sky position. We sample randomly over inclination $|\cos i|$. We calculate the extinction using mwdust. The positions at which Gaia BH1 and BH2 were detected are marked as the red stars. For Gaia BH1, the detectability as a function of sky position clearly shows the pattern of the scanning law. For Gaia BH2, the detectability does not have a dependence on the sky position. Bottom panel: Probability of inclusion of Gaia BH1 (left) and Gaia BH2 (right) in Gaia DR3 NSS astrometric orbits catalog as a function of the system’s distance. We fix all parameters as listed in Table 2, and only vary the distance of the system. We sample randomly over inclination $|\cos i|$. We calculate the extinction using mwdust. The distances at which Gaia BH1 and Gaia BH2 were detected at are denoted by the dotted black lines.
@@ 5.3 Inferences on the population of Galactic detached black holes
For Gaia BH1-like sources, we assume a $1M_{\odot}$ star with absolute magnitude $M_{G,0}=+4.6$. We consider orbital periods $200010<P_{orb}<$ 2000 days, spanning roughly the range of detected NSS astrometric orbits. We will refer to these as Sun-like star + BH binaries, or Sun+BH binaries for short.
For Gaia BH2-like sources, we assume a $1M_{\odot}$ star with absolute magnitude $M_{G,0}=+1.8$. We consider orbital periods $2000300<P_{orb}<$ 2000 days. The minimum orbital period is 300 days such that the semimajor axis of the orbit is larger than the typical red giant radius. We will refer to these as red giant + BH binaries, or RG+BH binaries for short.
@@
Figure 10 shows the orbital periods, parallaxes, and black hole masses for the full population vs. detected systems. On average, 3.5% of Sun+BH binaries with orbital periods $200010<P_{orb}<$ 2000 days within 2 kpc are Gaia DR3 detectable, and 16% of RG+BH binaries with orbital periods $2000300<P_{orb}<$ 2000 days within 2 kpc are Gaia DR3 detectable [^10] On average, 7.0% of Sun+BH binaries with orbital periods $2000300<P_{orb}<$ 2000 days within 2 kpc are Gaia DR3 detectable.. The period distribution unsurprisingly looks similar to the NSS astrometric orbits catalog. Sun+BH systems are generally difficult to detect beyond 1 kpc. RG+BH systems can be detected to at least 2 kpc because the red giant is much more luminous than a Sun-like star despite having the same mass. There is no strong dependence of the detectability of the system on black hole mass. The size of the photocenter orbit semimajor axis $a_{0}$ (Equation 1) increases by a factor of only 2.4 for a $5M_{\odot}$ vs. $50M_{\odot}$ black hole and $1M_{\odot}$ star, even though the black hole mass increases by a factor of 10.
@@
We can use this model to make a simple argument on the occurrence rate of detached black holes in the Milky Way in various orbital period ranges. We first consider Sun+BH binaries with orbital periods $400100<P_{orb}<$ 400 days, $1000400<P_{orb}<$ 1000 days, and $20001000<P_{orb}<$ 2000 days, and separately RG+BH binaries with orbital periods $1000400<P_{orb}<$ 1000 days and $20001000<P_{orb}<$ 2000 days. We assume that Gaia BH1 (Sun+BH with an orbital period $400100<P_{orb}<$ 400 days) and BH2 (RG+BH with an orbital period $20001000<P_{orb}<$ 2000 days) are the only two star + BH binaries in the Gaia DR3 astrometric orbits catalog. We then use our model to calculate the number of detectable detached BHs within 1 kpc of the Sun within these period ranges. For the period ranges where no BH has been detected, we take this to be the upper limit on the number of those BH binaries. Assuming there are $\approx$100-200 billion stars in the Galaxy, 160 million, or 0.1% of those stars are within 1 kpc of the Sun; to convert our number of objects to within 1 kpc to the number of objects in the Galaxy, we multiply by 1000.
Table 3 lists the detection probabilities. We find that the Galaxy hosts $5000\pm 2000$ Sun+BH binaries with $400100<P_{orb}<$ 400 days, and $20,000\pm 4000$ RG+BH binaries with $20001000<P_{orb}<$ 2000 days. We also find that there are at most 2000 Sun+BH binaries with $1000400<P_{orb}<$ 1000 day, 20,000 Sun+BH binaries with $20001000<P_{orb}<$ 2000 day, and 20,000 RG+BH binaries with $20001000<P_{orb}<$ 2000 day. Note the limits and constraints are weaker for the long-period end, due to the sharp decrease in number of solutions beyond 1000 days.
@@
Although the previous paragraphs treats the constraints on Sun+BHs separately from RG+BHs separately, in reality these constraints are related, as Sun-like stars will eventually evolve into red giants. Because the evolutionary timescale of the red giant phase is only a $\sim 10\%$ of a star’s main sequence lifetime, the number of RG+BHs should also be a few percent of the number of Sun+BH binaries in a given period range. This is in tension with the results for the $20001000<P_{orb}<$ 2000 range, as our model predicts more RG+BH binaries than Sun+BH binaries. This is because the empirical filter that removes long-period solutions probabilistically places an upper limit on the number of solutions. For sources that pass the $\texttt{ruwe}\ >1.4$ and $\varpi/\sigma_{\varpi}>20,000\textrm{ days}/P_{orb}$ filters, this limit is $\sim 4\%$ for $20001000<P_{orb}<$ 2000. Because of this, the number of RG+BHs in Table 3 is probably an overestimate.
@@
$P_{orb}$ range (day) Prob. detect N($d<1$ kpc)
@@
$1000400<P_{orb}<$ 1000 0.50 $<2$
$20001000<P_{orb}<$ 2000 0.05 $20\pm 4$
Note. — Number of black holes within 1 kpc of the Sun. Upper limits and uncertainties are $1\sigma$. To convert these results to the number of black holes in the entire Milky Way, multiply the values in the N($d<1$ kpc) column by 1000.
@@ 6.1 Discrepancy in eccentricity
Figure 11: Top panel: eccentricity distribution correction factor derived from the forward modeling. Bottom panel: NSS eccentricity distribution for Sun-like stars with orbital periods $1000100<P_{orb}<$ 1000 days, before and after correction. Note that we only plot orbits with $e>0.01$. The blue line shows a fit to the corrected distribution: a beta function with $\alpha=1.7$, $\beta=3.2$, rescaled with a minimum of 0.017 and width of 0.98. The ruwe  filter does not have a strong bias against eccentricity, so the raw and corrected distributions look very similar.
There is a serious discrepancy between the Gaia NSS astrometric orbit and forward modeled eccentricity distribution, with the forward model predicting significantly more eccentric binaries than are present in the Gaia astrometric orbits catalog. We hypothesize that the starting eccentricity distribution used in the forward model is incorrect and leads to this discrepancy.
@@
We use our forward model to determine a correction that can be applied to the observed NSS eccentricity distribution to convert it into the true underlying eccentricity distribution. We calculate the ratio of the normalized distribution of mock NSS astrometric orbits solution to the normalized distribution of the birth population as a function of eccentricity in four period bins. The reason for binning the results by period is that the eccentricity distribution is a function of period in the forward model (Equation 9). We then average the ratios across the period bins, and fit a quadratic through this average to determine a correction factor on the eccentricity distribution as a function of eccentricity (Figure 11). As expected, the astrometric cascade has a slight preference for filtering out high-eccentricity systems. This is due to the $\texttt{ruwe}>1.4$ filter (§4.1.1).
El-Badry et al. (2024) find a similar discrepancy between the Gaia NSS astrometric orbit and their forward modeled eccentricity distribution, with their forward model also predicting more eccentric binaries than are present in the Gaia astrometric orbits catalog. However, their discrepancy is not as significant as found in this work; they find their astrometric cascade produces a significantly stronger bias against the detection of eccentric orbits than we do. This bias is due to the eccentric orbits having larger astrometric uncertainties $\sigma_{a_{0}}$ because they are preferentially sampled near apastron; these solutions are then removed by the $a_{0}/\sigma_{a_{0}}$ filter. Although we do model the $a_{0}/\sigma_{a_{0}}$ filter in this work, our model for $\sigma_{a_{0}}$ is only a function of the number of visibility periods $n_{vis}$ and magnitude $G$ and does not account for the eccentricity or period of the orbit (§4.2). Thus, it does not capture the behavior observed in El-Badry et al. (2024), and the correction derived in Figure 11 thus only corrects the bias due to to the ruwe  filter.
We apply our correction to the eccentricity distribution in the NSS astrometric orbits catalog for Sun-like stars (which we roughly define as sources with $64<M_{G,0}<$ 6) and orbital periods $1000100<P_{orb}<$ 1000 (Figure 11) [^11] We note that that these two requirements do not actually change the shape of the eccentricity distribution as compared to the full NSS catalog, they just reduce the number of sources.. We find that the corrected distribution is very similar to the raw distribution (i.e., in our model, the detection probability for a binary is largely independent of eccentricity), and that the corrected Gaia NSS astrometric orbits catalog still contains significantly fewer high eccentricity binaries than predicted by Moe & Di Stefano (2017).
@@
6.2 Limitations of the approach
The filters described in §4.2 and 4.3 are empirical, and we do not attempt to derive the relations from first principles nor from the scanning law of the spacecraft. The empirical period filters attributed to the loss of orbital solutions assigned to acceleration solutions in the model cascade likely cannot be derived from first principles. In addition, it does not include potential correlations with other parameters (e.g., eccentricity) which might affect the probability of a source being given an acceleration solution. The empirical period filter is likely adequate for population studies, but less ideal when trying to infer properties of one single system. In addition, the empirical period filters do not map directly to a step in the model cascade, as the ruwe  and $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$ filters do. To do this would require forward modeling individual observations. For a more detailed approach that actually does simulate individual Gaia observations including the scanning law, see El-Badry et al. (2024) (further discussion is also presented in §6.5).

### v2

@@ 5.3 Probability of finding Gaia black holes
Note. — Properties of Gaia BH1 and BH2. * subscript denotes properties of the stellar companion, $\bullet$ subscript denotes properties of the black hole. All values except $M_{G,*}$ and $A_{G}$ are taken from Nagarajan et al. (2024) and El-Badry et al. (2023b) for Gaia BH1 and BH2, respectively. We note for Gaia BH2 that our extinction map suggests much less extinction ($\approx 0.4$ mag) than the map used by El-Badry et al. (2023b) and so our inferred extinction-corrected absolute magnitude of the star $M_{G,0*}$ in Gaia BH2 is fainter. We do not calculate the probability for Gaia BH3 Gaia Collaboration et al. (2024) because its orbital period is longer than 1500 days.
@@
We apply our forward modeling framework to determine the probability of inclusion of Gaia BH1 (El-Badry et al., 2023a; Chakrabarti et al., 2023) and Gaia BH2 (El-Badry et al., 2023b) in the Gaia DR3 NSS astrometric orbits catalog. We do not calculate the probability for Gaia BH3 (Gaia Collaboration et al., 2024) because its orbital period is much longer than 1500 days, where our ruwe  model breaks down. We calculate the probability as a function of sky position, orbital period, eccentricity, parallax, and the component magnitudes and masses. We integrate over random inclinations that are uniform in $|\cos i|$.
@@
We find the probability of Gaia BH1 ((RA, Dec) = (262.17, –0.58) deg, $P_{orb}=186$ day, $e=0.45$, $\varpi=2.09$ mas, $M_{*}=0.93M_{\odot}$, $M_{\bullet}=9.62M_{\odot}$, $G_{*}=13.77$ mag) being included in the Gaia DR3 NSS astrometric orbit catalog to be 38%. We also vary the location of BH1 on the sky and its distance from the Sun (Figure 11) and consider how its probability of inclusion in the Gaia DR3 NSS astrometric orbit catalog changes. Since Gaia BH1 is located near the ecliptic plane, it had a relatively low probability of inclusion; toward the ecliptic poles, the inclusion probability would have been closer to 60–80%. Gaia BH1 is also relatively close by, at a distance of about 500 pc. Had it been at the same (RA, Dec) but slightly farther away, at 600 pc or more, it would not have made it into the NSS catalog. Conversely, had it been slightly closer, 400 pc or less, it would have had $>80\%$ probability of making it into the catalog. This suggests that if there are Gaia BH1-like systems within a few hundred pc of the Sun, they should have been included in the Gaia DR3 NSS astrometric orbits catalog; the lack of such detections suggests that there are no or few Gaia BH1-like sources within 300–400 pc of the Sun.
@@
We find the probability of Gaia BH2 ((RA, Dec) = (207.57, –59.24), $P_{orb}=1277$ day, $e=0.52$, $\varpi=0.86$ mas, $M_{*}=1.07M_{\odot}$, $M_{\bullet}=8.94M_{\odot}$, $G_{*}=12.28$ mag) being included in the Gaia DR3 NSS astrometric orbit catalog to be 27%. We also vary the location of BH2 on the sky and its distance from the Sun (Figure 11) and consider how its probability of inclusion in the Gaia DR3 NSS astrometric orbit catalog varies. Since Gaia BH2 has a long orbital period, the main factor of whether it made it into the NSS catalog is whether or not it was assigned an acceleration solution or not. The inclusion probability shows a weak dependence on ecliptic latitude, and does not vary significantly ($<10$%) across the sky. Gaia BH2 is around 1.2 kpc from the Sun. Had it been at the same (RA, Dec) but beyond 3 kpc, it would not have been included in the NSS catalog. Conversely, had it been closer (between 500 and 1000 pc), it would have a somewhat higher probability ($50-60$%) of making it into the NSS catalog.
The results here for BH1 are consistent with the results in El-Badry et al. (2024), which find a 39% probability of inclusion in the Gaia DR3 NSS astrometric orbit catalog. The probability of inclusion for BH1 as a function of distance (Figure 15 of El-Badry et al. 2024) is also very similar. For BH2, El-Badry et al. (2024) finds a higher probability of 48%. The probability of inclusion for BH2 as a function of distance (Figure 15 of El-Badry et al. 2024) is also similar, albeit their inclusion probabilities are generally about $10-20$% higher. Since our model was tuned to match that of El-Badry et al. (2024), we might expect the results to be more similar. This may be due to the probabilistic nature of the acceleration cascade modeling; $P_{orb}$ and $a_{0}/\sigma_{AL}$ may not capture the full dependencies of probability acceptance, leading to large scatter. This suggests the methodology may be reliable only to a factor of ${\sim}2$ beyond $P_{orb}\gtrsim 500$ days.
@@
Figure 11: Top panels: Probability of inclusion of Gaia BH1 (left) and Gaia BH2 (right) in Gaia DR3 NSS astrometric orbits catalog as a function of sky position (192 equally spaced points on sky); the color of the circle denotes the probability of inclusion. Note that the color bar scales in the two panels are different. We fix all parameters as listed in Table 2, and only vary the sky position. We sample randomly over inclination $|\cos i|$. We calculate the extinction using mwdust. The positions at which Gaia BH1 and BH2 were detected are marked as the red stars. For Gaia BH1, the detectability as a function of sky position clearly shows the pattern of the scanning law. The ecliptic plane receives fewer scans than the ecliptic poles (Lindegren et al., 2012), reducing the astrometric precision and in turn the size of a detectable astrometric orbit. Gaia BH1 thus tends to be more reliably detected around the ecliptic poles than the ecliptic plane. For Gaia BH2, the detectability does not have as strong a dependence on the sky position. Bottom panel: Probability of inclusion of Gaia BH1 (left) and Gaia BH2 (right) in Gaia DR3 NSS astrometric orbits catalog as a function of the system’s distance. We fix all parameters as listed in Table 2, and only vary the distance of the system. We sample randomly over inclination $|\cos i|$. We calculate the extinction using mwdust. The distances at which Gaia BH1 and Gaia BH2 were detected at are denoted by the dotted black lines.
@@ 5.4 Inferences on the population of Galactic detached black holes
For Gaia BH1-like sources, we assume a $1M_{\odot}$ star with absolute magnitude $M_{G,0}=+4.6$. We consider orbital periods $150030<P_{orb}<$ 1500 days, spanning roughly the range of detected NSS astrometric orbits. We will refer to these as Sun-like star + BH binaries, or Sun+BH binaries for short.
For Gaia BH2-like sources, we assume a $1M_{\odot}$ star with absolute magnitude $M_{G,0}=+1.8$. We consider orbital periods $1500300<P_{orb}<$ 1500 days. The minimum orbital period is 300 days such that the semimajor axis of the orbit is larger than the typical red giant radius. We will refer to these as red giant + BH binaries, or RG+BH binaries for short.
@@
Figure 12 shows the orbital periods, parallaxes, and black hole masses for the full population vs. detected systems. On average, 5% of Sun+BH binaries with orbital periods $150030<P_{orb}<$ 1500 days within 2 kpc are Gaia DR3 detectable, and 27% of RG+BH binaries with orbital periods $1500300<P_{orb}<$ 1500 days within 2 kpc are Gaia DR3 detectable [^14] On average, 10% of Sun+BH binaries with orbital periods $1500300<P_{orb}<$ 1500 days within 2 kpc are Gaia DR3 detectable.. The period distribution unsurprisingly looks similar to the NSS astrometric orbits catalog. Sun+BH systems are generally difficult to detect beyond 1 kpc. RG+BH systems can be detected to at least 2 kpc because the red giant is much more luminous than a Sun-like star despite having the same mass. There is no strong dependence of the detectability of the system on black hole mass within the $5-80M_{\odot}$ range. The size of the photocenter orbit semimajor axis $a_{0}$ (Equation 1) increases by a factor of only 2.4 for a $5M_{\odot}$ vs. $50M_{\odot}$ black hole and $1M_{\odot}$ star, even though the black hole mass increases by a factor of 10.
@@
We can use this model to make a simple argument on the occurrence rate of detached black holes in the Milky Way in various orbital period ranges. We first consider Sun+BH binaries with orbital periods $400100<P_{orb}<$ 400 days, $1000400<P_{orb}<$ 1000 days, and $15001000<P_{orb}<$ 1500 days, and separately RG+BH binaries with orbital periods $1000400<P_{orb}<$ 1000 days and $15001000<P_{orb}<$ 1500 days. We assume that Gaia BH1 (Sun+BH with an orbital period $400100<P_{orb}<$ 400 days) and BH2 (RG+BH with an orbital period $15001000<P_{orb}<$ 1500 days) are the only two star + BH binaries in the Gaia DR3 astrometric orbits catalog, as the studies that have identified candidates in the astrometric NSS catalog (Andrews et al., 2022; Shahaf et al., 2023) have been followed up and not yielded detections. Additional candidates identified in the Gaia BH1 discovery papers (El-Badry et al., 2023a; Chakrabarti et al., 2023) also have been ruled out.
We then use our model to calculate the number of detectable detached BHs within 2 kpc of the Sun within these period ranges. For the period ranges where no BH has been detected, we take this to be the upper limit on the number of those BH binaries. Assuming there are 100 billion stars in the Galaxy, 1 billion, or 1% of those stars are within 2 kpc of the Sun; to convert our number of objects within 2 kpc to the number of objects in the Galaxy, we multiply by 100.
Table 3 lists the detection probabilities. We find that the Galaxy hosts $3000\pm 500$ Sun+BH binaries with $400100<P_{orb}<$ 400 days, and $400\pm 200$ RG+BH binaries with $15001000<P_{orb}<$ 1500 days. We also find that there are $<800$ Sun+BH binaries with $1000400<P_{orb}<$ 1000 day, $<12,000$ Sun+BH binaries with $15001000<P_{orb}<$ 1500 day, and $<300$ RG+BH binaries with $1000400<P_{orb}<$ 1000 day. These uncertainties are assuming Poisson statistics; they do not account for systematic errors (e.g., errors in the forward model, or the assumed number of stars in the Galaxy.)
@@
Although the previous paragraphs treats the constraints on Sun+BHs separately from RG+BHs separately, in reality these constraints are related, as Sun-like stars will eventually evolve into red giants. Because the evolutionary timescale of the red giant phase is only a $\sim 10\%$ of a star’s main sequence lifetime, the number of RG+BHs should also be a few percent of the number of Sun+BH binaries in a given period range. This is generally consistent with the results we find. For the $1000400<P_{orb}<$ 1000 day range, we only have upper limits for RG+BH and Sun+BH binaries and so this is not particularly constraining. For the $15001000<P_{orb}<$ 1500 day range, this implies that the number of RG+BH binaries are likely on the lower side of the estimated range, and for the Sun+BH binaries, the number of systems will be close to the upper limit and they just barely managed to escape inclusion in the NSS astrometric orbits catalog.
@@
$P_{orb}$ range (day) Prob. detect N($d<2$ kpc)
@@
$1000400<P_{orb}<$ 1000 0.36 $<3$
$15001000<P_{orb}<$ 1500 0.25 $4\pm 2$
Note. — Number of black holes within 2 kpc of the Sun. Upper limits and uncertainties are $1\sigma$, assuming Poisson errors. To convert these results to the number of black holes in the entire Milky Way, multiply the values in the N($d<2$ kpc) column by 100.
@@ 6.1 Discrepancy in eccentricity
There is a discrepancy between the Gaia NSS astrometric orbit and forward modeled eccentricity distribution, with the forward model predicting significantly more eccentric binaries than are present in the Gaia astrometric orbits catalog. We hypothesize that the starting eccentricity distribution used in the forward model is incorrect and leads to this discrepancy.
@@
Figure 13: Forward model assuming different birth eccentricity distributions. The birth distribution of eccentricities for orbital periods within $150030<P_{orb}<$ 1500 days is shown in the dotted blue line. The forward modeled eccentricity distribution is shown in the solid blue line. The distribution of eccentricities in the NSS astrometric orbits catalog is shown in black. The default prescription in Moe & Di Stefano (2017) (left panel) produces an excess of high-eccentricity orbits. Assuming a flat eccentricity distribution (left middle panel) $\eta=0$ is a slight improvement. Assuming a distribution with $\eta=-0.3$ (right middle panel) matches the long-eccentricity tail well, but overestimates the number of circular orbits. A Rayleigh distribution with $\sigma_{e}=0.33$ (right panel) provides the best match to the observed eccentricity distribution.
As a test of our forward model to see if a different eccentricity distribution would result in a better match to the observed NSS distribution, we re-run the population synthesis, replacing Equation 8 from Moe & Di Stefano (2017) with $\eta=0$ and $-0.3$. All other parts of the population synthesis remain the same. We then re-run the forward model and look at the resulting eccentricity distribution.
Figure 13 compares the different eccentricity prescriptions. Although neither of them provides a great match, they are improvements at the high eccentricity end. The issue with a power law parametrization is that any decrease in high-eccentricity solutions is compensated by more circular solutions. Therefore, a different functional form is likely needed. The observed NSS eccentricity distribution peaks at $e\approx 0.2-0.3$. Qualitatively, the birth distribution needs to have a concentration of solutions around $e=0.2-0.3$ to match the observed Gaia NSS distribution.
Wu et al. (2025) propose a Rayleigh distribution
$p(e)=\frac{e}{\sigma_{e}^{2}}\exp\left[-\frac{e^{2}}{2\sigma_{e}^{2}}\right]$ (9)
to model the eccentricity distribution of Gaia DR3 NSS sources. They find that for their “gold sample”, which consists of $\sim 3000$ Sun-like binaries with $10^{2}<P_{orb}<10^{3}$ days within 150 pc, the best-fit scale parameter is $\sigma_{e}=0.336$, which includes a completeness correction to account for NSS selection effects. We test this using our forward model. For our original synthetic population, for any source with $\log_{10}P_{orb}>0.7$, we instead draw their eccentricities from Equation 9 with $\sigma_{e}=0.33$. Because the Rayleigh distribution has support over $[0,\infty)$, we remove any sources that have $e>1$. Following Moe & Di Stefano (2017), we also removed sources where the eccentricity exceeded
$e_{max}(P_{orb})=1-\left(\frac{P_{orb}}{2\textrm{ d}}\right)^{-2/3}$ (10)
to ensure that the components have Roche lobe fill factors $\lesssim 70\%$ at periastron. We simply enforce a hard cutoff at $e_{max}$; in contrast, COSMIC applies a linear turnover in the number of eccentric solutions at $e/e_{max}=0.8$, such that the eccentricity distribution is continuous at $e/e_{max}=0.8$ and zero at $e=e_{max}$. Finally, we again run our NSS forward model to determine the resulting eccentricity distribution (rightmost panel of Figure 13), which now closely reproduces the observed NSS distribution. As noted in Wu et al. (2025), the birth and NSS forward modeled eccentricity distributions are very similar, meaning the eccentricity distribution is approximately complete. This is because a Rayleigh distribution with $\sigma_{e}\approx 0.3$ has very few high-eccentricity solutions, where the distribution is most affected by selection effects; it is not generally true of any eccentricity distribution.
@@
6.2 Strengths and limitations of the approach
In the spirit of keeping calculations fast and analytic in this work, we construct empirical probabilistic functions to approximate the fitting of epoch 1-D astrometry. This has both strengths and limitations, namely the tradeoff between speed and simplicity vs. computational expense and accuracy.
The main strength of a fast simulation is for exploring parameter space. It requires approximately $10-20$ seconds to run the full astrometric cascade of El-Badry et al. (2024) for each source that receives an orbital solution, or about 5000 CPU hours to produce a full mock DR3 binary catalog. In contrast, it requires approximately 2 seconds to run 100,000 evaluations of the empirical forward model of this work at a fixed position on sky; most of that time comes from querying the 3-D dust map. Of course, it is important to note that the empirical corrections to the analytic prescription are dependent on having the expensive epoch modeling simulations already in hand. It is analogous to significant upfront costs in training machine learning models, but later having substantially lower costs when using the model to make predictions. Finally, the simple approach also allows for a more intuitive understanding of each step in the cascade and elucidates the complex selection function.
The speed and simplification comes at the expense of detail and potentially accuracy. The major limitation of our analytic models is that they are limited to certain regions of parameter space. Most notably, as discussed earlier, our model does not cover orbital periods $P_{orb}>1500$ days. In addition, when trying to understand one specific system in detail, individual epoch modeling will likely be more accurate and can provide specific insight. One way to take advantage of the strengths of both the fast and detailed approaches are to use the fast simulations to identify potentially interesting regions of parameter space, then to use the detailed simulations to validate and refine those results.